A survey is taken of males in a large state University using a SRS. 72% of the 100 males reported that college was easier than high school. Create a 95% confidence interval for the true percentage of males at the University who feel college is easier than high school.
1
Expert's answer
2019-11-12T09:56:55-0500

"Standard\,error=\sqrt{\frac{p×(1-p)}{n}}=\sqrt{\frac{0.72×0.28}{100}}=0.045"

"z-value\,for\,95\%\,confidence=1.96"

"Upper\,bound=p+z×SE=0.72+1.96×0.045=0.808"

"Lower\,bound=p-z×SE=0.72-1.96×0.045=0.632"

Answer: 72%±4.5%